# Nearly Optimal Sample Complexity of Offline KL-Regularized Contextual Bandits under Single-Policy Concentrability

### v1

@@ Abstract
KL-regularized policy optimization has become a workhorse in learning-based decision making, while its theoretical understanding is still very limited. Although recent progress has been made towards settling the sample complexity of KL-regularized contextual bandits, existing sample complexity bounds are either $\widetilde{O}(\epsilon^{-2})$ under single-policy concentrability or $\widetilde{O}(\epsilon^{-1})$ under all-policy concentrability. In this paper, we propose the first algorithm with $\widetilde{O}(\epsilon^{-1})$ sample complexity under single-policy concentrability for offline contextual bandits. Our algorithm is designed for general function approximation and based on the principle of pessimism in the face of uncertainty. The core of our proof leverages the strong convexity of the KL regularization, and the conditional non-negativity of the gap between the true reward and its pessimistic estimator to refine a mean-value-type risk upper bound to its extreme. This in turn leads to a novel covariance-based analysis, effectively bypassing the need for uniform control over the discrepancy between any two functions in the function class. The near-optimality of our algorithm is demonstrated by an $\widetilde{\Omega}(\epsilon^{-1})$ lower bound. Furthermore, we extend our algorithm to contextual dueling bandits and achieve a similar nearly optimal sample complexity.
@@ 1 Introduction
Due to the data-hungry and instable nature of reinforcement learning (RL), divergences that are straightforward to estimate via Monte Carlo methods or amenable to constrained optimization stand out from numerous candidates (Rényi, 1961; Müller, 1997; Basseville, 2013) as regularizers; the former family of which is typically $f$-divergence because any of them is an expectation under one policy (thus easy for on-policy estimation) (Levine, 2018; Levine et al., 2020), and the latter are naturally cases of Bregman divergence (Bregman, 1967) induced by strongly convex functions because they are suitable for scalable optimization algorithms over the probability simplex by design (Lattimore and Szepesvári, 2020; Lattimore, 2024). In particular, Kullback-Leibler (KL) divergence is the only one at the intersection of Bregman divergence and $f$-divergence (Jiao et al., 2014, Theorem 5), indicating that it is uniquely favorable among common choices from both computational and statistical aspects. Therefore, the KL-regularized RL objective is arguably the most popular one in practice:
$J(\pi)=\mathbb{E}_{\pi}[r]-\frac{1}{\eta}\mathsf{KL}({\pi}\|{\pi^{\mathsf{ref}}}),$ (1.1)
@@
where $r$ is a reward function, $\pi$ is the policy to be optimized, $\pi^{\mathsf{ref}}$ is a reference policy, $\mathsf{KL}({\pi}\|{\pi^{\mathsf{ref}}})$ is the reverse KL divergence, and $\eta>0$ is the inverse temperature. When $\pi^{\mathsf{ref}}$ is chosen to be the uniform distribution, the objective function reduces to the (Shannon) entropy-regularized RL objective that encourages diverse actions and enhances robustness (Williams, 1992; Ziebart et al., 2008; Ziebart, 2010; Levine and Koltun, 2013; Fox et al., 2015; Levine et al., 2016; Haarnoja et al., 2018; Richemond et al., 2024; Liu et al., 2024). The KL-regularized objective has also been widely used in the RL fine-tuning of large language models (LLMs) (Ouyang et al., 2022; Rafailov et al., 2023, 2024), where $\pi^{\mathsf{ref}}$ is a pretrained LLM.
There has also been a surge of interest in understanding the principle behind KL-regularized RL. Ahmed et al. (2019); Liu et al. (2019) studied by ablation the effect of entropy regularization on the stability and policy improvement in policy optimization, the regret of which has been rigorously settled under the classic online mirror descent framework (Cai et al., 2020; He et al., 2022; Ji et al., 2023). Neu et al. (2017) unified popular KL-regularized policy optimization algorithms under a convex optimization framework, but the interplay of which with data were left untouched. A series of work (Geist et al., 2019; Vieillard et al., 2020; Kozuno et al., 2022) then analyzed the sample complexity of algorithms using KL w.r.t. the previous iteration or/and entropy regularizers with improved dependence on the effective horizon in discounted Markov decision processes (MDPs). However, the metric of suboptimality in all these studies is still with respect to the unregularized reward maximization objective, and its optimal sample complexity is $O(\epsilon^{-2})$ for finding the $\epsilon$-optimal policy.
Several recent studies (Xiong et al., 2024; Xie et al., 2024; Zhao et al., 2024) switched the focus to suboptimality guarantee with respect to the regularized objective in (1.1). In particular, Xiong et al. (2024) proposed an Offline GSHF algorithm via the principle of pessimism in the face of uncertainty, and proved $O(\epsilon^{-2})$ sample complexity under single-policy concentrability (See Section 2.1 for detailed definitions of concentrability.). On the other hand, the sharp analysis in Zhao et al. (2024) yields the optimal sample complexity $O(\epsilon^{-1})$, but requires all-policy concentrability (Zhao et al., 2024, Definition 2.6), i.e., the data-generating policy $\pi^{\mathsf{ref}}$ is required to cover the entire function class for all possible policies. Xie et al. (2024) studied token-level Markov decision processes (MDPs) and proposed a KL-regularized RL algorithm named XPO, which achieves $O(\epsilon^{-2})$ sample complexity under all-policy concentrability. In fact, even for offline contextual bandits (Foster and Rakhlin, 2023, Chapter 3) with $\mathrm{i.i.d.}$ data, which is among the simplest setting for offline decision making, existing sample complexity under KL-regularization are still primitive. Thus, a natural question arises:
Is the $\widetilde{O}(\epsilon^{-1})$ sample complexity w.r.t. 1.1 achievable for offline KL-regularized contextual bandits under single-policy concentrability?
We answer this question affirmatively by proposing a novel yet minimalist approach based on the principle of pessimism (Jin et al., 2021). Our algorithm can deal with general reward function which belongs to a reward function class with finite covering entropy. The optimal sample complexity is also due to our proof technique, which leverages the strong convexity of KL regularization and the conditional non-negativity of the gap between the true reward and its pessimistic estimator to refine a mean-value-type risk upper bound to its extreme. This in turn leads to a novel covariance-based analysis, effectively bypassing the need for uniform control over the discrepancy between any two functions in the function class.
Table 1: Comparison of sample complexity for finding $\epsilon$-optimal policy for offline contextual bandits. $\mathrm{polylog}(\epsilon^{-1})$ and the dependence of other parameters are omitted here. C(D)Bs stands for contextual (dueling) bandits. The last column indicates whether the bounds hold under single-policy concentrability.
Reference [^1] The four existing upper bounds are adapted from the implicit form in Xiong et al. (2024, Theorem 3.1) and Zhao et al. (2024, Theorem 3.3 and Theorem 4.4.). See Section 2.2 for details. CBs CDBs Single?
Xiong et al. (2024) $\widetilde{O}(\epsilon^{-2})$ $\widetilde{O}(\epsilon^{-2})$
Zhao et al. (2024) $\widetilde{O}(\epsilon^{-1})$ $\widetilde{O}(\epsilon^{-1})$
This work $\widetilde{O}(\epsilon^{-1})$ $\widetilde{O}(\epsilon^{-1})$
Lower bounds $\widetilde{\Omega}(\epsilon^{-1})$ $\widetilde{\Omega}(\epsilon^{-1})$  [^2] The lower bounds in terms of $\epsilon$ are the same under both single- and all-policy concentrability, which are detailed in Remarks 2.11 and 3.8 respectively.
@@ 1.1 Contributions
  • We propose a new algorithm dubbed KL-PCB and develop a novel proof technique that simultaneously utilizes the curvature of KL-regularized objectives and integrates pessimism with a novel covariance-based argument, for deriving its $\widetilde{O}(\epsilon^{-1})$ sample complexity under single-policy concentrability.
  • To demonstrate the versatility of our techniques beyond absolute reward feedback, we also extend our results to the offline contextual dueling bandit (CDB) setting, and provide a tight sample complexity analysis.
  • The matching lower bounds (up to $\mathrm{polylog}$ factors) in both settings suggest that our sample complexity results are nearly optimal, which renders a complete picture of the sample complexity for KL-regularized offline contextual under single policy concentrability.
For the ease of comparison, we summarize the relevant sample complexity results along with the corresponding coverage assumptions in Table 1.
@@ 1.2 Additional Related Work
However, in terms of risk decomposition, to the best of our knowledge, none of these pessimism-based analyses really goes beyond the performance difference lemma (Foster and Rakhlin, 2023, Lemma 13) or simulation lemma (Foster and Rakhlin, 2023, Lemma 23); both of which are not able to capture the strong concavity of KL-regularized objectives even in the bandit setting. The algorithmic idea of using pessimistic least-square estimators under general function approximation in Jin et al. (2021); Di et al. (2023) is similar to ours, but their suboptimality gap is bounded by the sum of bonuses, the direct adaptation of which to our objective cannot lead to the desired sample complexity.
Offline Contextual Dueling Bandits
@@
Notation
Boldfaced lower case (resp. upper case) letters are reserved for vectors (resp. matrices). The sets ${\mathcal{S}}$ and $\mathcal{A}$ are assumed to be countable throughout this paper to avoid measure-theoretic arguments. For nonnegative sequences $\{x_{n}\}$ and $\{y_{n}\}$, we write $x_{n}=O(y_{n})$ if $\limsup_{n\to\infty}{x_{n}}/{y_{n}}<\infty$, and write $y_{n}=\Omega(x_{n})$ if $x_{n}=O(y_{n})$. We further employ $\widetilde{O}(\cdot)$ and $\widetilde{\Omega}(\cdot)$ to hide $\mathrm{polylog}$ factors. Given a positive definite $\bm{\Sigma}\in\mathbb{R}^{d\times d}$ and $\mathbf{x}\in\mathbb{R}^{d}$, we denote the vector’s Euclidean norm by $\|\mathbf{x}\|_{2}$ and define $\|\mathbf{x}\|_{\bm{\Sigma}}=\sqrt{\mathbf{x}^{\top}\bm{\Sigma}\mathbf{x}}$. For coutable $\mathcal{X}$ and $\mathcal{Y}$, we denote the family of probability kernels from $\mathcal{X}$ to $\mathcal{Y}$ by $\Delta(\mathcal{Y}|\mathcal{X})$. For a pair of probability measures $P\ll Q$ on the same space, the KL divergence from $P$ to $Q$ is ${\mathsf{KL}}\left(P\|Q\right)\coloneqq\int\log({{\,\mathrm{d}}P}/{{\,\mathrm{d}}Q}){\,\mathrm{d}}P$, and their total variation (TV) distance is ${\mathsf{TV}}\left(P\|Q\right)\coloneqq 0.5\int|{\,\mathrm{d}}P-{\,\mathrm{d}}Q|$. For $f:\mathcal{X}\to\mathbb{R}$, its infinity norm is denoted by $\|f\|_{\infty}\coloneqq\sup_{x\in\mathcal{X}}|f(x)|$.
@@ 2.1 Problem Setup
We consider contextual bandit, which is denoted by a tuple $({\mathcal{S}},\mathcal{A},r,\pi^{\mathsf{ref}})$. Specifically, ${\mathcal{S}}$ is the context space, $\mathcal{A}$ is the action space and $r:{\mathcal{S}}\times\mathcal{A}\to[0,1]$ is the reward function. In the offline setting, the agent only has access to an $\mathrm{i.i.d.}$ dataset $\mathcal{D}=\{(s_{i},a_{i},r_{i})\}_{i=1}^{n}$. Here $s_{i}\in{\mathcal{S}}$ is the state sampled from a distribution $\rho$, $a_{i}\in\mathcal{A}$ is the action taken from a behavior policy, and $r_{i}$ is the observed reward given by $r_{i}=r(s_{i},a_{i})+\varepsilon_{i}$, where $\varepsilon_{t}$ is $1$-sub-Gaussian (Lattimore and Szepesvári, 2020, Definition 5.2). In this work, we consider the KL-regularized objective
@@
where $\pi^{\mathsf{ref}}$ is a known reference policy and the “inverse temperature” $\eta$ controls the intensity of regularization. For simplicity, we assume that $\pi^{\mathsf{ref}}$ is also the behavior policy that generates the dataset $\mathcal{D}$. This type of “behavior regularization” has been also been studied in Zhan et al. (2022). The unique optimal policy $\pi^{*}\coloneqq\mathop{\mathrm{argmax}}_{\pi\in\Delta(\mathcal{A}|{\mathcal{S}})}J(\pi)$ is in the following closed-form (See, e.g., Zhang 2023, Proposition 7.16)
$\pi^{*}(\cdot|s)\propto\exp\big{(}\eta\cdot r(s,\cdot)\big{)},\forall s\in{\mathcal{S}}.$ (2.2)
A policy $\pi$ is said to be $\epsilon$-optimal if $J(\pi^{*})-J(\pi)\leq\epsilon$ and the goal of the agent is to find one such policy based on $\mathcal{D}$. To ensure $\epsilon$-optimality is achievable, we assume that $r$ lies in a known function class $\mathcal{F}\subset({\mathcal{S}}\times\mathcal{A}\to[0,1])$, from which the agent obtains her estimator $\widehat{f}$.
Assumption 2.1 (General Function Approximation).
There exists a known function class $\mathcal{F}:{\mathcal{S}}\times\mathcal{A}\to[0,1]$ such that $\exists\ f^{*}\in\mathcal{F}$ such that $f^{*}=r$.
We also need a standard condition to control the complexity of $\mathcal{F}$ through the notion of covering number (Wainwright, 2019, Definition 5.1).
@@
For any $\epsilon_{c}>0$, the $\epsilon_{c}$-covering number $\mathcal{N}_{\mathcal{F}}(\epsilon_{c})$ of $\mathcal{F}$ is $\mathrm{poly}(\epsilon_{c}^{-1})$.

### v2

# Towards a Sharp Analysis of Offline Policy Learning for $f$-Divergence-Regularized Contextual Bandits

@@ Abstract
Although many popular reinforcement learning algorithms are underpinned by $f$-divergence regularization, their sample complexity with respect to the regularized objective still lacks a tight characterization. In this paper, we analyze $f$-divergence-regularized offline policy learning. For reverse Kullback–Leibler (KL) divergence, arguably the most commonly used one, we give the first $\widetilde{O}(\epsilon^{-1})$ sample complexity under single-policy concentrability for contextual bandits, surpassing existing $\widetilde{O}(\epsilon^{-1})$ bound under all-policy concentrability and $\widetilde{O}(\epsilon^{-2})$ bound under single-policy concentrability. Our analysis for general function approximation leverages the principle of pessimism in the face of uncertainty to refine a mean-value-type argument to its extreme. This in turn leads to a novel moment-based technique, effectively bypassing the need for uniform control over the discrepancy between any two functions in the function class. We further propose a lower bound, demonstrating that a multiplicative dependency on single-policy concentrability is necessary to maximally exploit the strong convexity of reverse KL. In addition, for $f$-divergences with strongly convex $f$, to which reverse KL does not belong, we show that the sharp sample complexity $\widetilde{\Theta}(\epsilon^{-1})$ is achievable even without single-policy concentrability. In this case, the algorithm design can get rid of pessimistic estimators. We further extend our analysis to dueling bandits, and we believe these results take a significant step toward a comprehensive understanding of $f$-divergence-regularized policy learning.
@@ 1 Introduction
Due to the data-hungry and instable nature of reinforcement learning (RL), divergences that are straightforward to estimate via Monte Carlo or amenable to constrained optimization stand out from numerous candidates (Rényi, 1961; Csiszár, 1967; Müller, 1997; Basseville, 2013) as regularizers; the former family is typically $f$-divergence (Rényi, 1961) because any of them is an expectation, for which empirical average is a good proxy (Levine, 2018; Levine et al., 2020); and the latter class subsumes those with nice positive curvatures (e.g., Bregman divergence (Bregman, 1967) induced by strongly convex functions). In particular, Kullback-Leibler (KL) divergence is the only one at the intersection of $f$-divergence and Bregman divergence (Jiao et al., 2014, Theorem 5), indicating its theoretical advantage among common choices from both computational and statistical aspects. Also, the KL-regularized RL objective is arguably the most popular one in practice:
$J(\pi)=\mathbb{E}_{\pi}[r]-\eta^{-1}\mathsf{KL}({\pi}\|{\pi^{\mathsf{ref}}}),$ (1.1)
@@
where $r$ is the reward, $\pi^{\mathsf{ref}}$ is a reference policy, $\mathsf{KL}({\pi}\|{\pi^{\mathsf{ref}}})$ is the reverse KL divergence, and $\eta>0$ is the inverse temperature. When $\pi^{\mathsf{ref}}$ is uniform, 1.1 reduces to the entropy-regularized objective that encourages diverse actions and enhances robustness (Williams, 1992; Ziebart et al., 2008; Levine and Koltun, 2013; Levine et al., 2016; Haarnoja et al., 2018; Richemond et al., 2024; Liu et al., 2024). KL regularization has also been widely used in the RL fine-tuning of large language models (Ouyang et al., 2022; Rafailov et al., 2023), where $\pi^{\mathsf{ref}}$ is the base model. Given its widespread use, there has been a surge of interest in understanding the role of KL regularization in RL by both empirical studies (Ahmed et al., 2019; Liu et al., 2019) and theoretical analysis (Geist et al., 2019; Vieillard et al., 2020; Kozuno et al., 2022). There are also lines of research on KL regularization in online learning (Cai et al., 2020; He et al., 2022; Ji et al., 2023) and convex optimization (Neu et al., 2017). However, the performance metric in most of these studies is still the unregularized reward maximization objective, under which the sample complexity is at least $\Omega(\epsilon^{-2})$. (See Section A.1 for detailed discussions)
Table 1: Comparison of sample complexity bounds for finding $\epsilon$-optimal policy for offline contextual bandits with KL- and (strongly convex) $f$-divergence regularization. Constants and $\mathrm{polylog}$ factors are omitted here except the metric entropy $\log\mathcal{N}$. “Reverse-KL” stands for KL-regularized contextual bandits and “$f$-divergence w/ s.c., $f$” for the counterpart with an $\alpha$-strongly convex $f$. The two existing upper bounds are adapted from the implicit form in Xiong et al. (2024, Theorem 3.1) and Zhao et al. (2024, Theorem 3.3 and Theorem 4.4.), of which the detailed adaptions are deferred to Appendix A. Also see Remark 2.8 for discussion on the relationship between $D^{2}_{\pi^{*}}$ and $C^{\pi^{*}}$
Regularizer Xiong et al. (2024) Zhao et al. (2024) This work
Reverse KL Upper $d\epsilon^{-2}$ $\eta D^{2}\epsilon^{-1}\log\mathcal{N}$ $\eta D^{2}_{\pi^{*}}\epsilon^{-1}\log\mathcal{N}$
Lower - $\eta\epsilon^{-1}\log\mathcal{N}$ $\eta C^{\pi^{*}}\epsilon^{-1}\log\mathcal{N}$
$f$-divergence w/ s.c. $f$ Upper - - $\alpha^{-1}\eta\epsilon^{-1}\log\mathcal{N}$
Lower - - $\alpha^{-1}\eta\epsilon^{-1}\log\mathcal{N}$
Several recent papers (Xiong et al., 2024; Xie et al., 2024; Zhao et al., 2024; Foster et al., 2025) switched the focus to analyzing the suboptimality defined via the regularized objective (1.1), under which an $\Omega(\epsilon^{-1})$ sample complexity is possible (Zhao et al., 2024, Theorem 3.6). However, even restricted to the pure $\mathrm{i.i.d.}$ setting, existing analyses in this vein either result in still $\widetilde{O}(\epsilon^{-2})$ bounds (Xiong et al., 2024; Xie et al., 2024) or has a stringent all-policy concentrabiity dependency in their upper bounds (Zhao et al., 2024) (See Section 2.1 for details on coverage conditions). Thus, there are by far no perfectly matching bounds down to both the dependency of $\epsilon^{-1}$ and concrete coverage conditions for KL-regularized offline decision making. Moreover, these theoretical analyses set KL as the right target by default; but reverse KL is the $f$-divergence with $f(x)=x\log x$, which is merely convex. Therefore, it is also unknown whether those choices with even nicer (e.g., strongly convex) $f$, whose performance against the reward maximization objective can be provably promising (Zhan et al., 2022; Gabbianelli et al., 2024; Huang et al., 2024), can enjoy a better coverage dependency in the sample complexity under the corresponding regularized objectives. These perspectives motivate an open problem:
What is the weakest coverage condition for offline policy learning to be minimax optimal with respect to $f$-divergence-regularized objectives?
We are the first to concretely investigate this question by proposing sharp sample complexity analyses for two representative ad subclasses of $f$-divergence. In particular, for contextual bandits with KL regularization, we present sample complexity guarantee with linearly dependence on single-policy coverage ratio. Our lower bound further indicates that a multiplicative dependency on single-policy concentrability is necessary. Surprisingly, for $f$-divergence with $\alpha$-strongly-convex $f$, we prove nearly matching sample complexity bounds of $\widetilde{\Theta}(\alpha^{-1}\eta\epsilon^{-1})$ for the first time, settling the dependence on coverage completely. For the ease of comparison, we adapt previous sample complexity bounds to our setting and notation, and summarize them in Table 1.
@@ 1.1 Contributions
  • For arguably the most commonly used KL regularization, we are the first to employ the principle of pessimism (Jin et al., 2021) to achieve a nearly tight sample complexity under single-policy concentrability. Our analysis roadmap leverages the strong convexity of KL and pessimism of the reward estimator, to refine a mean-value-type risk upper bound to its extreme which in turn leads to a novel moment-based analysis, effectively bypassing the need for uniform control over the discrepancy between any two functions in the function class.
  • Under KL-regularization, we prove a lower bound growing linearly the density-ratio-based single-policy concentrability, which also strictly improved upon previous works (Zhao et al., 2024; Foster et al., 2025) in offline setting. This lower bound manifests the necessity of a multiplicative dependency on single-policy concentrability and firstly closes the gap between upper and lower bounds under a technically nontrivial case.
  • For any $f$-divergence with strongly convex $f$, we are able to design a truly lightweight algorithm free of pessimism-based gadgets and still obtain the $\widetilde{\Theta}(\epsilon^{-1})$ sample complexity certified by a matching lower bound without coverage conditions.
  • We extend all algorithmic ideas and hard instance constructions above to $f$-divergence-regularized contextual dueling bandits, achieving $\widetilde{\Theta}(\epsilon^{-1})$ sample complexity bounds. Moreover, all algorithms work for general reward function classes with finite metric entropy.
@@ 1.2 Additional Related Work
However, in terms of risk decomposition, to the best of our knowledge, none of these pessimism-based analyses really goes beyond the performance difference lemma (Foster and Rakhlin, 2023, Lemma 13) or simulation lemma (Foster and Rakhlin, 2023, Lemma 23); both of which are not able to capture the strong concavity of KL-regularized objectives even in the bandit setting. The algorithmic idea of using pessimistic least-square estimators under general function approximation in Jin et al. (2021); Di et al. (2023) is similar to ours, but their suboptimality gap is bounded by the sum of bonuses, which cannot lead to the desired sample complexity of our objective by direct adaption.
Offline contextual dueling bandits.
@@
Notation.
The sets ${\mathcal{S}}$ and $\mathcal{A}$ are assumed to be countable throughout the paper. For nonnegative sequences $\{x_{n}\}$ and $\{y_{n}\}$, we write $x_{n}=O(y_{n})$ if $\limsup_{n\to\infty}{x_{n}}/{y_{n}}<\infty$, $y_{n}=\Omega(x_{n})$ if $x_{n}=O(y_{n})$, and $y_{n}=\Theta(x_{n})$ if $x_{n}=O(y_{n})$ and $x_{n}=\Omega(y_{n})$. We further employ $\widetilde{O}(\cdot),\widetilde{\Omega}(\cdot)$, and $\widetilde{\Theta}$ to hide $\mathrm{polylog}$ factors. For countable $\mathcal{X}$ and $\mathcal{Y}$, we denote the family of probability kernels from $\mathcal{X}$ to $\mathcal{Y}$ by $\Delta(\mathcal{Y}|\mathcal{X})$. For $g:\mathcal{X}\to\mathbb{R}$, its infinity norm is denoted by $\|g\|_{\infty}\coloneqq\sup_{x\in\mathcal{X}}|g(x)|$. For a pair of probability measures $P\ll Q$ on the same space and function $f:\mathbb{R}_{+}\to\mathbb{R}$, their $f$-divergence is $D_{f}(P\|Q)\coloneqq\int f({\,\mathrm{d}}P/{\,\mathrm{d}}Q){\,\mathrm{d}}Q$. Specifically, when $f(x)=x\log x$, $f$-divergence becomes KL divergence denoted as ${\mathsf{KL}}\left(P\|Q\right)\coloneqq\int\log({{\,\mathrm{d}}P}/{{\,\mathrm{d}}Q}){\,\mathrm{d}}P$, and when $f(x)=|x-1|/2$, it becomes the total variation (TV) distance, which is denoted as ${\mathsf{TV}}\left(P\|Q\right)\coloneqq 0.5\int|{\,\mathrm{d}}P-{\,\mathrm{d}}Q|$.
@@ 2.1 Problem Setup
We consider contextual bandit, which is denoted by a tuple $({\mathcal{S}},\mathcal{A},r,\pi^{\mathsf{ref}})$. Specifically, ${\mathcal{S}}$ is the context space, $\mathcal{A}$ is the action space and $r:{\mathcal{S}}\times\mathcal{A}\to[0,1]$ is the reward function. In the offline setting, the agent only has access to an $\mathrm{i.i.d.}$ dataset $\mathcal{D}=\{(s_{i},a_{i},r_{i})\}_{i=1}^{n}$. Here $s_{i}^{\prime}s$ are states sampled from $\rho\in\Delta({\mathcal{S}})$, $a_{i}\in\mathcal{A}$ is the action taken from a behavior policy, and $r_{i}$ is the observed reward given by $r_{i}=r(s_{i},a_{i})+\varepsilon_{i}$, where $\varepsilon_{t}$ is $1$-sub-Gaussian (Lattimore and Szepesvári, 2020, Definition 5.2). In this work, we consider the KL-regularized objective
@@
where $\pi^{\mathsf{ref}}$ is a known reference policy and the “inverse temperature” $\eta$ controls the intensity of regularization. For simplicity, we assume that $\pi^{\mathsf{ref}}$ is also the behavior policy that generates the dataset $\mathcal{D}$. This type of “behavior regularization” has been also been studied in Zhan et al. (2022). The unique optimal policy $\pi^{*}\coloneqq\mathop{\mathrm{argmax}}_{\pi\in\Delta(\mathcal{A}|{\mathcal{S}})}J(\pi)$ is given by (See, e.g., Zhang 2023, Proposition 7.16)
$\pi^{*}(\cdot|s)\propto\pi^{\mathsf{ref}}(\cdot|s)\exp\big{(}\eta\cdot r(s,\cdot)\big{)},\forall s\in{\mathcal{S}}.$ (2.2)
A policy $\pi$ is said to be $\epsilon$-optimal if $J(\pi^{*})-J(\pi)\leq\epsilon$ and the goal of the agent is to find one such policy based on $\mathcal{D}$. To ensure $\epsilon$-optimality is achievable, we assume that $r$ lies in a known function class ${\mathcal{G}}\subset({\mathcal{S}}\times\mathcal{A}\to[0,1])$, from which the agent obtains the estimator $\widehat{r}$. More specifically, we work with general function approximation under realizability, which is as follows.
Assumption 2.1.
For this known function class ${\mathcal{G}}\subset({\mathcal{S}}\times\mathcal{A}\to[0,1])$, $\exists{g^{*}}\in{\mathcal{G}}$ with ${g^{*}}=r$.
We also need a standard condition to control the complexity of ${\mathcal{G}}$ through the notion of covering number (Wainwright, 2019, Definition 5.1).
@@
For any $\epsilon_{c}>0$, the $\epsilon_{c}$-covering number $\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})$ of ${\mathcal{G}}$ is $\mathrm{poly}(\epsilon_{c}^{-1})$.
When ${\mathcal{G}}$ is the class of linear functions of dimension $d$ and radius $R$, the covering number of ${\mathcal{G}}$ is given by $\mathcal{N}_{{\mathcal{G}}}=O((1+R\epsilon^{-1})^{d})$ (Jin et al., 2020, Lemma D.6), indicating that the Assumption 2.3 on $\mathcal{N}_{{\mathcal{G}}}$ is mild.
Concentrability.
The data quality of $\mathcal{D}$ collected by $\pi^{\mathsf{ref}}$ is typically characterized by concentrability in offline RL (Farahmand et al., 2010; Chen and Jiang, 2019; Jiang and Xie, 2024), which quantifies the ability of the behavioral policy to generate diverse actions. We first define the density-ratio-based concentrability as follows.
Definition 2.4 (Density-ratio-based concentrability).
For policy class $\Pi$, reference policy $\pi^{\mathsf{ref}}$, the density-ratio-based all-policy concentrability $C^{\Pi}$ is $C^{\Pi}\coloneqq\sup_{\pi\in\Pi,s\in{\mathcal{S}},a\in\mathcal{A}}{\pi(a|s)}/{\pi^{\mathsf{ref}}(a|s)}$, whose single-policy counterpart under the optimal policy $\pi^{*}$ is $C^{\pi^{*}}\coloneqq\sup_{s\in{\mathcal{S}},a\in\mathcal{A}}{\pi^{*}(a|s)}/{\pi^{\mathsf{ref}}(a|s)}$.
In the definition above, small all-policy concentrability intuitively corresponds to $\text{supp}(\pi^{\mathsf{ref}})$ covering all possible inputs. On the other hand, small single-policy concentrability means that $\text{supp}(\pi^{\mathsf{ref}})$ only subsumes $\text{supp}(\pi^{*})$. In this paper, in addition to density-ratio-based concentrability, we also adopt the following $D^{2}$-based concentrabilites to better capturing the nature of function class ${\mathcal{G}}$. In detail, we first introduce $D^{2}$-divergence as follows.
Definition 2.5.
Given a function class ${\mathcal{G}}\subset({\mathcal{S}}\times\mathcal{A}\to\mathbb{R})$ and a fixed policy $\pi$, define the $D^{2}$-divergence $D^{2}_{{\mathcal{G}}}((s,a);\pi)$ as
$\sup_{g,h\in{\mathcal{G}}}\frac{\big{(}g(s,a)-h(s,a)\big{)}^{2}}{\mathbb{E}_{(s^{\prime},a^{\prime})\sim\rho\times\pi}[(g(s^{\prime},a^{\prime})-h(s^{\prime},a^{\prime}))^{2}]}.$
The “eluder dimension”-type Definition 2.5 is directly inspired by Di et al. (2023); Zhao et al. (2024), the intuition behind which is that given $(s,a)\in{\mathcal{S}}\times\mathcal{A}$, a small $D^{2}$-divergence indicates that for two functions $g$ and $h$, if they are close under the behavior policy $\pi$, then they will also be close on such pair $(s,a)$. Therefore, the $D^{2}$-divergence quantifies how well the estimation on dataset collected by the behavior policy $\pi$ can be generalized to a specific state-action pair. We are now ready to define the two notions of concentrability conditions.
Assumption 2.6 (All-policy concentrability).
Given a reference policy $\pi^{\mathsf{ref}}$, there exists $D<\infty$ such that $D^{2}=\sup_{(s,a)\in{\mathcal{S}}\times\mathcal{A}}D^{2}_{{\mathcal{G}}}((s,a);\pi^{\mathsf{ref}})$.
Assumption 2.6 indicates that the errors on any state-action pairs can be bounded by the error on the samples from $\rho\times\pi$ up to a factor $D$, whose relaxed counterpart under the same $\pi^{\mathsf{ref}}$ is as follows.
Assumption 2.7 (Single-policy concentrability).
$D_{\pi^{*}}^{2}\coloneqq\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}D^{2}_{{\mathcal{G}}}((s,a);\pi^{\mathsf{ref}})<\infty$.
Assumption 2.7 indicates that the errors on the distributions of state-action pairs $\rho\times\pi^{*}$ can be bounded by the error on the samples from $\rho\times\pi^{\mathsf{ref}}$ up to some constant. For both types, the single-policy concentrability assumption is strictly weaker than the all-policy concentrability assumption. However, in general, the two quantities characterizing single-policy concentrability $C^{\pi^{*}}$ and $D^{2}_{\pi^{*}}$ cannot be bounded by each other up to constant factors.